OpenStudy (anonymous):

findP(E or F) if P(E and F)= 0.15

jimthompson5910 (jim_thompson5910):

Hint: P(E or F) = P(E) + P(F) - P(E and F)

OpenStudy (anonymous):

P(E)= 0.25 P(F)= 0.45

jimthompson5910 (jim_thompson5910):

plug in the given info into that equation, then evaluate

OpenStudy (anonymous):

then what do I do with 0.15?

OpenStudy (anonymous):

\[P(E\cap F)=P(E)+P(F)-P(E\cup F)\\ 0.15=0.25+0.45-P(E\cup F)\\ P(E\cup F)=\cdots\]